Can someone please help me find the circumference of a circle with a radius of 45cm?

Thanks

Answer by hailu(13)

C=Πd
C=3.14*90
C=282.6
OR
C=2Πr
C=2*3.14*45
C=6.28*45
C=282.6

make sure u memorize those two formulas,it'll help u solve other problems. i'm hoping u know that diameter is twice it's radius,thats how i got the answer.
happy holidays.